1.

Colter Steel has $4,700,000 in assets. Temporary current assets Permanent current assets Fixed assets Total assets $ 1,400,000 1,520,000 1780,000 $ 4,700,000 Assume the term structure of interest rates becomes inverted, with short-term rates going to 11 percent and long-term rates 5 percentage points lower than short-term rates. Earnings before interest and taxes are $1,000,000. The tax rate is 20 percent. Earnings after taxes

2.

Carmens Beauty Salon has estimated monthly financing requirements for the next six months as follows January February March $ 8,300 April 2,300 May 3,300 June $8,300 9,300 4,300 Short-term financing will be utilized for the next six months. Projected annual interest rates are January February March 7.0% April 8.0% May 11.0% June 14.0% 12.0% 12.0% a. Compute total dollar interest payments for the six months. (Round your monthly interest rate to 2 decimal places when expressed as a percent. Round your interest payments to the nearest whole cent.) Total dollar interest payments b-1. Compute the total dollar interest payments if long-term financing at 12 percent had been utilized throughout the six months? (Round your monthly interest rate to 2 decimal places when expressed as a percent. Round your interest payments to the nearest whole cent.) Total dollar interest payments
Answer #1

Answer (1). Long Term Financing equals Permanent Current Assets and Fixed Assets. Short term financing equals temporary current assets.

Permanent Current Assets = $1,520,000

Fixed Assets = $1,780,000

Total Long Term Finance raised

to support long-term assets = --------------------------

$ 3,300,000

--------------------------

Interest on long term finance = (11-5)% * 3300000 = $198,000

Short term financing = $1,400,000

Interest on short term finance = 11% * 1400000 = $ $154,000

Total Interest = $(198000+154000)

= $352,000

Now, Earnings before interest and tax = $ 1,000,000

Interest = ($352,000)

Earnings Before Tax = $648,000

Tax (0.2*648000)= ($129,600)

Earnings After TAx = $ 518,400

----------------------------------------------------------------------------------------------------

(2) (a)

Month Amount Annual Interest Rate Monthly Interest Rate Monthly Interest
January 8300 7% 0.58% 48.14
February 2300 8% 0.67% 15.41
March 3300 11% 0.92% 30.36
April 8300 14% 1.17% 97.11
May 9300 12% 1% 93
June 4300 12% 1% 43
Total

327.02

(2) (b)

Month Amount Annual Interest Rate Monthly Interest Rate Monthly Interest
January 8300 12% 1% 83
February 2300 12% 1% 23
March 3300 12% 1% 33
April 8300 12% 1% 83
May 9300 12% 1% 93
June 4300 12% 1% 43
Total 358